find the mean of the probability distribution of the random variable X and interpret its value, which can take only the values of 3,5, and 7, given that P (3) = ⁷/³⁰, P(5) = ⅓ and P(7) = ¹³/³⁰

1
Expert's answer
2022-02-21T18:01:36-0500

"\mu=3*\frac{7}{30}+5*\frac{1}{3}+7*\frac{13}{30}=\frac{162}{30}=5.4."

The average value of a random variable X over a large number of experiments equals 5.4.